E=a body freely falling from a height h describes 7h/16 in the last

Grade:12th pass

1 Answers

Vikas TU
7 years ago
First eqn . write as:
h = 0 + 0.5gt^2........................(1)
 
7h/16 = 0 + 0.5g(t – 1/2).............(2)
SOLVE both the eqns.
and get h value.